# Photon Spheres and shadow of Schwarzschild black hole on the EUP framework

### 04 Mar 2026 v1

@@ 1 Introduction
On 10 April 2019, the global astronomical community simultaneously released the first image of a black hole. The black hole is located at the centre of the nearby galaxy M87, which is situated 55 million light-years from Earth. The mass of this black hole is 6.5 billion times that of the Sun. As one of the most enigmatic and destructive celestial objects in the universe K190611238 ; K190611239 ; K190611240 ; K190611241 ; K190611242 ; K190611243 ; KL25 ; KL26 , the discovery of such supermassive black holes not only challenges traditional theories of black hole formation but also compels the scientific community to re-examine the fundamental nature and evolutionary mechanisms of black holes.
@@
In the vicinity of a black hole, gravitational forces exerted by the object result in the deflection of light rays, a phenomenon known as gravitational lensing H250820419 ; R220602161 ; S220612820 ; K9904193 ; P180100860 . This phenomenon leads to a substantial decrease in the observational intensity within the region delimited by the black hole, culminating in the formation of a low-luminosity area on the distant observation plane-the black hole shadow. The boundary of this shadow is defined by the trajectories of photons that can escape to infinity, and its geometric dimensions exceed those of the event horizon. It is noteworthy that the black hole shadow demonstrates a significant correlation with spacetime geometry; the morphology of the shadow exhibits distinct variations among diverse black hole types, thereby offering indispensable observational indicators for the investigation of black hole characteristics S230100113 . Consequently, shadow analysis has become an effective tool for estimating black hole parameters V210507101 ; Y251106219 ; P250904261 ; J251023047 ; W251106017 ; A139945 ; X250803020 ; K250208388 ; S230902709 ; M20248437 ; H202484350 ; R892782020 ; S201108023 ; M210311417 ; S220602488 and testing gravitational theories such as general relativity A210108261 ; Z211101752 ; R220700078 ; S220507787 .
@@
In principle, substantial advances have been made in the geometric characterisation of the photon sphere and the black hole shadow. In the context of static, spherically symmetric spacetimes, the photon sphere can be delineated by the conditions of zero geodesic curvature and zero Gaussian optical curvature. This provides a coordinate-independent mathematical framework for the study of zero-geodesic circular orbits. Any deviation from the standard Bekenstein-Hawking entropy-area relation (and consequently from Schwarzschild geometry) naturally leads to modifications of the spacetime metric. Building on this principle, A251213769 systematically investigates modified metrics corresponding to different forms of entropy modification, calculates the resulting photon sphere radii and shadow dimensions, and reveals that each entropy modification induces characteristic optical shift effects. This study establishes new constraint boundaries for various entropy deformation parameters by comparing with Event Horizon Telescope (EHT) observations of the galactic centre black hole $\text{Sgr}{{\text{A}}^{*}}$.
@@
Ref. H250820419 focuses on minimal deformation models for two types of Schwarzschild black holes: the Kazakov-Solodukhin (KS) black hole arising from quantum corrections, and the Ghosh-Kumar (GK) black hole derived from nonlinear electrodynamics (NED). It is evident that both of these approaches result in minimal deformation through straightforward algebraic substitutions of the radial coordinate in the metric function. Each approach introduces a ”hair” parameter, which serves to characterize the degree of deformation. The study found that, upon analysis, the parameter in question has a significant effect on the black hole shadow. In comparison to Schwarzschild black holes, the event horizon, photon sphere radius and critical collision parameter all increased for KS black holes, while the corresponding quantities decreased for GK black holes. When considered in conjunction with EHT observational data, this work establishes constraints on the permissible parameter ranges for both types of black hole models.
@@
The manifestation of quantum gravity corrections in the context of black hole thermodynamics has also attracted considerable attention. Examples of such corrections include the Bekenstein-Hawking entropy correction and the extended generalized uncertainty principle (EGUP) correction to the Hawking radiation temperature. Ref. H250815467 investigates the impact of quantum corrections on the thermodynamic properties of Reissner-Nordström black holes containing quintessence matter within a dark energy background. Ref. U260300916 undertakes a systematic investigation of the thermodynamic properties and PVP-V critical behaviour of RN-AdS black holes surrounded by perfect fluid dark matter (PFDM) within the EGUP framework. Ref. Y250814953 undertakes a unified topological and geometric analysis of charged AdS black holes embedded in quintessence fields, incorporating infrared gravitational corrections introduced by the extended uncertainty principle (EUP). Ref. A250610621 investigates the accelerated radiation phenomenon exhibited by a two-level atom in free fall around a GUP-modified Schwarzschild black hole. Ref. B240515558 calculates the EGUP corrections to the thermodynamic quantities of RN-AdS black holes, including the Hawking temperature, heat capacity, entropy function and pressure, under the influence of quintessence, a dark energy candidate with negative pressure.
@@
It is imperative to emphasise that any deviation from the Bekenstein-Hawking entropy-area relation will induce corresponding modifications to the spacetime metric. Recent studies A251104613 ; A251213769 demonstrate that, starting from a selected entropy functional, one can systematically derive the associated modified metric and its effective matter distribution, thereby establishing a direct entropy-geometry correspondence. Inspired by this and drawing on the approach in A250721663 that derives van der Waals black holes by equating the van der Waals system temperature with the Hawking radiation temperature, this paper investigates the modified metric corresponding to the Hawking radiation temperature with EUP corrections. The impact of the EUP effect on the position of the photon sphere and the shadow radius of black holes is analysed further, based on the derived modified metric. The results indicate that as the EUP parameter increases, the photon sphere radius expands relative to that of a Schwarzschild black hole, while the shadow radius contracts; the position of the black hole’s event horizon remains unchanged. This finding provides a novel perspective on the role of the uncertainty principle in shaping black hole shadow characteristics, with significant implications for deepening our cosmic understanding.
@@ 2 Heisenberg uncertainty principle
The expression for the EUP-modified Hawking temperature is given by H250815467 ; U260300916 ; Y250814953 ; A250610621 ; B240515558
@@
Using the Eqs. (2.1) and (2.3), the black hole entropy can be derived from the first law of thermodynamics, $dM=TdS$, which yields H250815467 ; U260300916 ; Y250814953 ; A250610621 ; B240515558 ,

### 08 Mar 2026 v2

@@ 1 Introduction
On 10 April 2019, the global astronomical community simultaneously released the first image of a black hole. The black hole is located at the centre of the nearby galaxy M87, which is situated 55 million light-years from Earth. The mass of this black hole is 6.5 billion times that of the Sun. As one of the most enigmatic and destructive celestial objects in the universe K190611238 ; K190611239 ; K190611243 ; KL25 ; KL26 , the discovery of such supermassive black holes not only challenges traditional theories of black hole formation but also compels the scientific community to re-examine the fundamental nature and evolutionary mechanisms of black holes.
@@
In the vicinity of a black hole, gravitational forces exerted by the object result in the deflection of light rays, a phenomenon known as gravitational lensing H250820419 ; R220602161 ; S220612820 ; K9904193 ; U260300916 ; P180100860 . This phenomenon leads to a substantial decrease in the observational intensity within the region delimited by the black hole, culminating in the formation of a low-luminosity area on the distant observation plane-the black hole shadow. The boundary of this shadow is defined by the trajectories of photons that can escape to infinity, and its geometric dimensions exceed those of the event horizon. It is noteworthy that the black hole shadow demonstrates a significant correlation with spacetime geometry; the morphology of the shadow exhibits distinct variations among diverse black hole types, thereby offering indispensable observational indicators for the investigation of black hole characteristics P241101261 . Consequently, shadow analysis has become an effective tool for estimating black hole parameters A250700455 ; S251207922 ; V947 ; S260401 ; P250904261 ; J251023047 ; W251106017 ; A139945 ; X250803020 ; K250208388 ; S230902709 ; M20248437 ; H202484350 ; R892782020 ; S201108023 ; M210311417 ; S220602488 and testing gravitational theories such as general relativity A210108261 ; Z211101752 ; V210507101 ; R220700078 ; S220507787 .
@@
In principle, substantial advances have been made in the geometric characterisation of the photon sphere and the black hole shadow C190412983 ; M240803241 ; E240909750 . In the context of static, spherically symmetric spacetimes, the photon sphere can be delineated by the conditions of zero geodesic curvature and zero Gaussian optical curvature. This provides a coordinate-independent mathematical framework for the study of zero-geodesic circular orbits. Any deviation from the standard Bekenstein-Hawking entropy-area relation (and consequently from Schwarzschild geometry) naturally leads to modifications of the spacetime metric. Building on this principle, A251213769 systematically investigates modified metrics corresponding to different forms of entropy modification, calculates the resulting photon sphere radii and shadow dimensions, and reveals that each entropy modification induces characteristic optical shift effects. This study establishes new constraint boundaries for various entropy deformation parameters by comparing with Event Horizon Telescope (EHT) observations of the galactic centre black hole $\text{Sgr}{{\text{A}}^{*}}$.
@@
Ref. H250820419 focuses on minimal deformation models for two types of Schwarzschild black holes: the Kazakov-Solodukhin (KS) black hole arising from quantum corrections, and the Ghosh-Kumar (GK) black hole derived from nonlinear electrodynamics (NED). It is evident that both of these approaches result in minimal deformation through straightforward algebraic substitutions of the radial coordinate in the metric function. Each approach introduces a ”hair” parameter, which serves to characterize the degree of deformation. The study found that, upon analysis, the parameter in question has a significant effect on the black hole shadow. In comparison to Schwarzschild black holes, the event horizon, photon sphere radius and critical collision parameter all increased for KS black holes, while the corresponding quantities decreased for GK black holes. When considered in conjunction with EHT observational data, this work establishes constraints on the permissible parameter ranges for both types of black hole models. Meanwhile, the Lorentzian-Euclidean black hole has recently been introduced as a geodesically complete spacetime, characterized by a signature change at the event horizon that prevents causal geodesics from reaching the central singularity at $r=0$. To identify deviations of this geometry from the standard Schwarzschild solution, the shadows produced by this black hole model have been systematically investigated in Refs. E260110806 ; S240417267 ; S250708431 .
@@
The manifestation of quantum gravity corrections in the context of black hole thermodynamics has also attracted considerable attention. Examples of such corrections include the Bekenstein-Hawking entropy correction and the extended generalized uncertainty principle (EGUP) correction to the Hawking radiation temperature. Ref. H250815467 investigates the impact of quantum corrections on the thermodynamic properties of Reissner-Nordström black holes containing quintessence matter within a dark energy background. Ref. B250908005 undertakes a systematic investigation of the thermodynamic properties and PVP-V critical behaviour of RN-AdS black holes surrounded by perfect fluid dark matter (PFDM) within the EGUP framework. Ref. Y250814953 undertakes a unified topological and geometric analysis of charged AdS black holes embedded in quintessence fields, incorporating infrared gravitational corrections introduced by the extended uncertainty principle (EUP). Ref. A250610621 investigates the accelerated radiation phenomenon exhibited by a two-level atom in free fall around a GUP-modified Schwarzschild black hole. Ref. B240515558 calculates the EGUP corrections to the thermodynamic quantities of RN-AdS black holes, including the Hawking temperature, heat capacity, entropy function and pressure, under the influence of quintessence, a dark energy candidate with negative pressure.
@@
It is imperative to emphasise that any deviation from the Bekenstein-Hawking entropy-area relation will induce corresponding modifications to the spacetime metric. Recent studies A251104613 demonstrate that, starting from a selected entropy functional, one can systematically derive the associated modified metric and its effective matter distribution, thereby establishing a direct entropy-geometry correspondence. Inspired by this and drawing on the approach in A250721663 that derives van der Waals black holes by equating the van der Waals system temperature with the Hawking radiation temperature, this paper investigates the modified metric corresponding to the Hawking radiation temperature with EUP corrections. The impact of the EUP effect on the position of the photon sphere and the shadow radius of black holes is analysed further, based on the derived modified metric. The results indicate that as the EUP parameter increases, the photon sphere radius expands relative to that of a Schwarzschild black hole, while the shadow radius contracts; the position of the black hole’s event horizon remains unchanged. This finding provides a novel perspective on the role of the uncertainty principle in shaping black hole shadow characteristics, with significant implications for deepening our cosmic understanding.
@@ 2 Heisenberg uncertainty principle
The expression for the EUP-modified Hawking temperature is given by H250815467 ; B250908005 ; Y250814953 ; A250610621 ; B240515558
@@
Using the Eqs. (2.1) and (2.3), the black hole entropy can be derived from the first law of thermodynamics, $dM=TdS$, which yields H250815467 ; B250908005 ; Y250814953 ; A250610621 ; B240515558 ,